OpenStudy (anonymous):

If f(x) = x + 4, find f(-7).

OpenStudy (anonymous):

f(-7) = -7 + 4 = -3

OpenStudy (anonymous):

f(x)=x+4 f(-7)=-7+4 = -3

OpenStudy (anonymous):

Which relation does not represent a function? Group A (-7, 3), (-3, 1), (1, -4), (5, -6) Group B (-4, 2), (1, 5), (-4, 0), (5, 6) Group C (2, 0), (4, 8), (6, 5), (10, 5) Group D: (-3, 6), (-1, 1), (0, 4), (3, 7)